# Exploring and Benchmarking the Planning Capabilities of Large Language Models

### v1

@@ 3.1 In-context learning
Logistics
The planning capabilities of GPT-4 and Gemini 1.5 models on the Logistics benchmark are shown in Figure 2(e) for PDDL and in Figure 2(b) for Natural Language. The 1-shot planning capability of Gemini 1.5 Pro reaches 43% for PDDL and for Natural Language 48%. Moreover for Gemini 1.5 Pro increasing the context consistently lead to better results, indicating that the model can make effective use of additional contexts. For Gemini 1.5 Flash and GPT-4 Turbo, the performance drops slight for PDDL and Natural Language.
Mini-Grid
Figure 2(f) and Figure 2(c) show the performance of GPT-4 Turbo and Gemini models as we increase the number of few-shot examples for PDDL and Natural Language, respectively. The Gemini models perform comparably for both PDDL and Natural Language, although GPT-4 Turbo appears to perform slightly better with Natural Language. Increasing the number of few-shot examples leads to better performance for all models. With 400 shots, specify model, Gemini 1.5 Pro reached 77% accuracy.
Trip Planning and Calendar Scheduling
Figure 3(a) shows the performance on Trip planning, Calendar Scheduling as we increase the number of few-shot examples, respectively. The 1-shot performance of the GPT-4 Turbo model seems to be better than the Gemini 1.5 Pro. We observe that, in both benchmarks, for both GPT-4 and Gemini 1.5 Flash the performance first increases with the number of shots and after a certain point, having more shots leads to worse model performance. However, for Gemini 1.5 Pro performance improves as the number of shot increases. Therefore, Gemini 1.5 Pro seems to be making more efficient use of additional shots compared to the other two. On the other hand GPT-4 performs better in one shot scenario compared to the other two models for Trip Planning, while Gemini 1.5 Pro has a higher accuracy in the 1-shot setting..
@@
Effect of better search strategies
Additionally, in Figure 4 we provide experimental evidence that methods such as MCTS and ToT can augment even Gemini 1.0 M GeminiTeam et al. (2024a) (which preceded Gemini 1.5 GeminiTeam et al. (2024b)), to be competitive with GPT-4 at smaller few-shot context lengths. Without these search procedures, Gemini 1.0 M fails to perform these planning tasks at any number of few-shot examples. Additional details and parameters for these search methods are included in Appendix B.3.
Figure 4: Calendar Scheduling task with search procedures (ToT and MCTS). We use Gemini 1.0 M to perform these procedures, and compare with GPT-4.
@@ 3.2 Supervised Fine Tuning (SFT)
In this section, we investigate the effect of Supervised Fine Tuning (SFT) with optimal plan on planning capability of LLMs. We specifically ran experiments on Gemini 1.0 S (GeminiTeam et al., 2024a) and investigate the effect of fine tuning on that on the planning performance for two different benchmarks with different levels of difficulty, namely, we look into 5 scenarios: BlocksWorld with 3-7 blocks, BlocksWorld with 8-9 blocks and BlocksWorld with 8-20 blocks, Logistics with 1-2 packets, Logistics with 3-5 packets. The data size and splits are documented in Appendix B.4.
The results are shown in Table 1. We observe that SFT leads to very high accuracy for some instances of both datasets and outperforms long context in-context learning. Moreover, the performance drops as the planning problem becomes more difficult.
@@ 3.3 Plan generalization
Here we investigate ability of LLMs to generalize their plans to unseen cases of different difficulty. For these experiments, we consider BlocksWorld and Logistics benchmarks with various difficulty levels used in Section 3.2. We look into performance of both SFT and ICL approaches. Tables 2 shows generalization performance for BlocksWorld in SFT setting for BlocksWorld 3-7, 8-20 split and Table 3 considers generalization between BlocksWorld 3-7 and 8-9 blocks split for both SFT and ICL with different models and number of shots. Table 4 depicts plan generalization for Logistics benchmark for splits of 1-2 and 3-5 packets.
Our analysis reveals several key findings: (1) Superiority of SFT: SFT consistently outperforms ICL across both benchmarks, even when utilizing a smaller model for SFT. This suggests that SFT’s explicit training process, focused on the specific task, leads to more effective learning and generalization. (2) In most ICL scenarios, training the model on easier instances first results in improved performance on harder examples, e.g., see Table 3, rows 1-4 and 6. (3) Limitations of hard example training: Contrary to some expectations, training the model exclusively on hard examples does not always translate to better performance on easier ones (for example, see Table 3 rows 2, 4-6). This suggests that a balanced approach, incorporating both easy and hard examples, might be optimal for achieving well-rounded performance.
@@
G 1.5 Flash (1 Shot) 26.4 / 13.7 32.9 / 12.3
G 1.5 Flash (70 Shot) 35.7 / 23.3 27.6 / 11.0
G 1.5 Pro (1 Shot) 40.0 / 25.6 39.0 / 20.3
G 1.5 Pro (70 Shot) 46.3 / 36.3 38.3 / 18.4
G 1.0 S (1 Shot) 3.68 / 0.331 2.99 / 1.35
G 1.0 S (70 Shot) 12.4 / 2.33 3.99 / 1.68
G 1.0 S (SFT) 96.3 / 81.6 96.0 / 92.6
Table 3: Testing BlocksWorld (PDDL representation) accuracy results by model on OOD tasks. For training we use in-context learning or fine-tuning. We consider a split with problems of size 3-7 blocks and 8-9 blocks. We use the shorthand G for Gemini.
@@
G 1.5 Flash (1 Shot) 18.3 / 1.35 26.7 / 1.00
G 1.5 Flash (30 Shot) 12.7 / 1.67 19.7 / 1.33
G 1.5 Pro (1 Shot) 35.3 / 9.03 57.6 / 7.01
G 1.5 Pro (30 Shot) 56.4 / 11.3 62.7 / 8.04
G 1.0 S (1 Shot) 7.0 / 0.0 5.33 / 0.0
G 1.0 S (30 Shot) 9.99 / 0.336 8.00 / 0.662
G 1.0 S (SFT) 99.8 / 10.8 98.0 / 63.4
Table 4: Testing Logistics (PDDL representation) accuracy results by model on OOD tasks. For training we use in-context learning or fine-tuning. We consider a split with problems of 1-2 packets and 3-5 packets, rest of parameters are the same in the splits. Due to space constraints we use the shorthand G for Gemini.
@@ 3.4 Comparison with Plan Bench
Val-BW GPT-4 1 NL 49.0 32.4 23.2
Val-BW G1.5P 1 NL 30.0 18.4 14.2
Val-BW G1.5P 1 PDDL 60.0 36.4 23.6
Val-BW G1.5P 2 PDDL 68.0 46.2 30.9
Our-BW G1.5P 1 NL 66.0 38.5 32.4
Our-BW G1.5P 2 NL 66.0 58.5 53.9
Our-BW G1.5P 1 PDDL 100 44.6 32.0
Our-BW G1.5P 2 PDDL 100 44.6 50.9
Table 5: Reporting accuracy (in %) to compare state of the art for different datasets and systems. Val-BW denotes BlocksWorld dataset as open-source by Valmeekam et al. (2023). We use GPT-4 to refer to GPT-4 Turbo, the shorthand G for Gemini, and P for Pro.
@@
We utilize the natural language prompts from Valmeekam et al. (2023) test them on Gemini 1.5 Pro. We observe that GPT-4 performs better with these prompts. For our dataset, no such difference is observed. Manual inspection reveals that especially 1-shot prompts need to be crafted carefully while few-shot or many-shot prompts are more robust. For instance, results improve when the prompts are more specific about the output format, changing from ’My plan is as follows’ to ’Your plan as plain text without formatting’, which enhances results for Gemini 1.5 Pro. Further, our prompts do not include explanations of the actions and we do not use color coding for the blocks but rather keep the names (e.g., blue block vs block b3).
4 Conclusion and Future Work
In this study, we comprehensively investigated the planning capabilities of Large Language Models (LLMs) through three key approaches: (1) Benchmark development (2) Learning and search techniques, (3) Generalization assessment. Our results demonstrate that SFT, ICL and search procedures all contribute significantly to enhancing the planning capabilities of LLMs.
In many-shot learning setting, comparing the performance of different LLMs reveals an intriguing phenomenon: while accuracy generally improves with increasing context, for some benchmarks, certain LLMs struggle to utilize additional context effectively, leading to a decline in performance. This observation holds implications for the future development and training of LLMs, potentially informing strategies to enhance their capacity to process and leverage extended contextual information.
Our investigation of plan generalization reveals three key findings: superiority of SFT, curriculum learning effectiveness and limitations of hard example training; suggesting that a balanced approach, incorporating both easy and hard examples, might be optimal for achieving well-rounded performance.
For future studies we are interested in two main aspects: Enabling plan generalization across diverse environments and Augmenting LLMs with replanning capabilities. These enhancements can unlock new levels of versatility and robustness in LLM-based planning systems, paving the way for their broader adoption in real-world applications.
5 Limitations
In this work we investigated a small number of large language models and did not cover every existing LLM such as Claude and Llama, in part due to compute and time constraints and in part due to some preliminary experiments where some of these models led to worse results compared to Gemini and GPT-4 for some of the approaches in the paper. For these reasons, and for the sake of environmental impact, we decided not to run extensive experiments on those models. Additionally, it is known that LLMs are sensitive to prompt engineering and it can be possible that our designed prompts were not necessarily optimal for any of the models we did include in this paper. In addition to prompt engineering there may also exist additional hyperparameters which can improve the performance of the methods described in this paper.
6 Ethics Statement
As is, our approach only teaches the LLM specific planning tasks mentioned in the paper which pose no significant ethical concerns. But we acknowledge that malicious actors might get inspiration from any work that teaches LLM a new capability and LLMs are prone to potential misuse and the improved planning ability can potentially be used for malicious intent. However, recommend careful consideration of all aspects of safety before relevant techniques are applied to the real world and suggest that the owners and engineers behind these models ensure that they filter out such requests.
The data used in our work is sourced from public datasets and our work complies with ACL Ethics Policy.
@@ B.2 Mappings PDDL to Natural Language
drive-truck param2 param3 param4 param5 Drive truck param2 from param3 to param4 in param5.
@@
unlock param2 param3 param4 param5 Unlock param2 at param3 using param4, which has param5.

### v2

@@ 3.1.1 Many-shot in-context learning
Logistics: The planning capabilities of GPT-4 and Gemini 1.5 models on the Logistics benchmark are shown in Figure 2(e) for PDDL and in Figure 2(b) for Natural Language. The 1-shot planning capability of Gemini 1.5 Pro reaches 43% for PDDL and for Natural Language 48%. Moreover for Gemini 1.5 Pro increasing the context consistently lead to better results, indicating that the model can make effective use of additional contexts. For Gemini 1.5 Flash and GPT-4, the performance drops slight for PDDL and Natural Language.
Mini-Grid: Figure 2(f) and Figure 2(c) show the performance of GPT-4 and Gemini models as we increase the number of few-shot examples for PDDL and Natural Language, respectively. The Gemini models perform comparably for both PDDL and Natural Language, although GPT-4 appears to perform slightly better with Natural Language. Increasing the number of few-shot examples leads to better performance for all models. With 400 shots, Gemini 1.5 Pro reaches 77% accuracy.
Trip Planning and Calendar Scheduling: Figure 3(a) shows the performance on the Trip Planning and Calendar Scheduling natural language tasks as we increase the number of few-shot examples, respectively. We observe that, in both benchmarks, for both GPT-4 and Gemini 1.5 Flash the performance first increases with the number of shots and after a certain point, having more shots leads to worse model performance. However, for Gemini 1.5 Pro performance improves as the number of shot increases. Therefore, Gemini 1.5 Pro seems to be making more efficient use of additional shots compared to the other two. On the other hand GPT-4 performs better in the 1-shot scenario compared to the other two models for Trip Planning, while Gemini 1.5 Pro has a higher accuracy in the 1-shot setting.
@@
3.1.2 Effect of inference time techniques
In addition to standard ICL, we consider inference time ICL methods that are based on constructing a chain-of-thought (Wei et al., 2022): Tree-of-Thought (ToT) (Yao et al., 2023), Monte Carlo Tree Search (MCTS) (Hao et al., 2023), and Debate-as-reasoning (Du et al., 2023).
In Figure 4 we provide experimental evidence that methods such as Debate-as-reasoning, MCTS, and ToT can augment Gemma2 27B (considerably smaller than Gemini 1.5 and GPT-4 models), to be competitive with GPT-4, Gemini 1.5 Pro and Gemini 1.5 Flash at smaller few-shot context lengths. Additional details and parameters for these search methods are included in Appendix B.3. These results demonstrate how inference-time reasoning procedures using smaller open-source models can perform competitively with larger foundational models in the natural language planning domain.
Figure 4: Calendar Scheduling (left) and Travel Scheduling (right) tasks with reasoning procedures (ToT, MCTS, and Debate). We use Gemma2 27B to perform these procedures, and compare with GPT-4 and Gemini 1.5 models.
However, the ability of these methods to scale to few-shot examples seems to be less significant than larger foundational models. Specifically, for the Travel Scheduling task, the larger models all outperform the CoT methods after 4 few-shot examples. It is unclear if this scaling trend is a result of the underlying model (Gemma2 27B) or the method itself.
Most interesting to note is the performance of debate-as-reasoning, which, despite not being an explicit search or planning strategy, works comparatively to MCTS in both tasks. This indicates that, for some planning tasks, even unstructured CoT style inference methods are competitive with explicit search-based CoT methods, and larger foundational models with ICL. This may indicate that, for natural language planning tasks, allowing the model to construct a CoT and consider multiple solutions may be more valuable then the specific method behind constructing the CoT.
@@ 3.2 Supervised Fine Tuning (SFT)
Supervised fine tuning (SFT) (Ouyang et al., 2022) has proved to be an effective method for teaching LLMs various capabilities. In this section, we investigate the effect of Supervised Fine Tuning (SFT) with optimal plan on planning capability of LLMs. We use the Fast-Downward classical planner to generate the optimal plan. We specifically ran experiments on
Gemini 1.0 S (GeminiTeam et al., 2024a) and investigate planning performance for two different benchmarks with different levels of difficulty, namely, we look into 5 scenarios: BlocksWorld with 3-7 blocks, 8-9 blocks or 8-20 blocks, and Logistics with 1-2 packets, or with 3-5 packets. The data size and splits are documented in Appendix B.4.
The results are shown in Table LABEL:tab:finetune_result. We observe that SFT leads to high accuracy for some instances of both datasets and outperforms many-shot ICL. The performance appears to drop as the planning problem becomes more difficult.
Table 1: Impact of SFT on accuracy, for BlocksWorld: instance of 3-7, 8-9 and 8-20 blocks, and for Logistics: instances of 1-2 and 3-5 packets.
@@ 3.3 Plan generalization
For any LLM application, the question of how well the method generalizes to out-of-training-distribution (OOD) inputs is always present. Here, we investigate how LLM planning generalizes. Plan generalization has three main categories: (1) Generalize to unseen instances of the same environment (2) Generalize to renaming of actions and objects (3) Generalize to unseen plan environments. We focus on the first category which sits at the core of the desired capability.
For these generalization experiments, we consider BlocksWorld and Logistics benchmarks with various difficulty levels (as described in Section 3.2). We look into performance of both SFT and ICL approaches. Tables LABEL:tab:plan-generalization-bw shows generalization performance for BlocksWorld in SFT setting for BlocksWorld 3-7, 8-20 split and Table 3 considers generalization between BlocksWorld 3-7 and 8-9 blocks split for both SFT and ICL with different models and number of shots. Table 4 depicts plan generalization for Logistics benchmark for splits of 1-2 and 3-5 packets.
Our analysis reveals several key findings: (1) Superiority of SFT: SFT consistently outperforms ICL across both benchmarks, even when utilizing a smaller model for SFT. This suggests that SFT’s explicit training process, focused on the specific task, leads to more effective learning and better generalization. (2) In most ICL scenarios, training the model on easier instances first results in improved performance on harder examples, e.g., see Table 3, rows 1-4 and 6. (3) Limitations of hard example training: Contrary to some expectations, training the model exclusively on hard examples does not always translate to better performance on easier ones (for example, see Table 3 rows 2, 4-6). This suggests that a balanced approach, incorporating both easy and hard examples, might be optimal for achieving well-rounded performance.
Table 3: OOD accuracy for BlocksWorld splits 3-7 and 8-9 blocks.
@@
Gemini 1.5 Flash (1 Shot) 26.4 / 13.7 32.9 / 12.3
Gemini 1.5 Flash (70 Shot) 35.7 / 23.3 27.6 / 11.0
Gemini 1.5 Pro (1 Shot) 40.0 / 25.6 39.0 / 20.3
Gemini 1.5 Pro (70 Shot) 46.3 / 36.3 38.3 / 18.4
Gemini 1.0 S (1 Shot) 3.68 / 0.331 2.99 / 1.35
Gemini 1.0 S (70 Shot) 12.4 / 2.33 3.99 / 1.68
Gemini 1.0 S (SFT) 96.3 / 81.6 96.0 / 92.6
Table 4: OOD accuracy for Logistics tasks splits of 1-2, 3-5 packets.
@@
Gemini 1.5 Flash (1 Shot) 18.3 / 1.35 26.7 / 1.00
Gemini 1.5 Flash (30 Shot) 12.7 / 1.67 19.7 / 1.33
Gemini 1.5 Pro (1 Shot) 35.3 / 9.03 57.6 / 7.01
Gemini 1.5 Pro (30 Shot) 56.4 / 11.3 62.7 / 8.04
Gemini 1.0 S (1 Shot) 7.0 / 0.0 5.33 / 0.0
Gemini 1.0 S (30 Shot) 9.99 / 0.336 8.00 / 0.662
Gemini 1.0 S (SFT) 99.8 / 10.8 98.0 / 63.4
@@ 3.4 Comparison with PlanBench
Table 5: Accuracy (in %) comparing the state of the art for different datasets and systems. Val-BW denotes BlocksWorld dataset as open-source by Valmeekam et al. (2023).
@@
Val-BW GPT-4 Turbo 1 NL 49.0 32.4 23.2
Val-BW Gemini 1.5 Pro 1 NL 30.0 18.4 14.2
Val-BW Gemini 1.5 Pro 1 PDDL 60.0 36.4 23.6
Val-BW Gemini 1.5 Pro 2 PDDL 68.0 46.2 30.9
Our-BW Gemini 1.5 Pro 1 NL 66.0 38.5 32.4
Our-BW Gemini 1.5 Pro 2 NL 66.0 58.5 53.9
Our-BW Gemini 1.5 Pro 1 PDDL 100 44.6 32.0
Our-BW Gemini 1.5 Pro 2 PDDL 100 44.6 50.9
@@
We utilize the natural language prompts from (Valmeekam et al., 2023) test them on Gemini 1.5 Pro. We observe that GPT-4 performs better with these prompts. For our dataset, no such difference is observed. Manual inspection reveals that especially 1-shot prompts need to be crafted carefully while few-shot or many-shot prompts are more robust. For instance, results improve when the prompts are more specific about the output format, changing from ’My plan is as follows’ to ’Your plan as plain text without formatting’, which enhances results for Gemini 1.5 Pro. Further, our prompts do not include explanations of the actions and we do not use color coding for the blocks but rather keep the names (e.g., blue block vs block b3).
4 Investigating Failure Cases
In this section, we analyze the outcome of various approaches to improving LLM planning performance, and dive into a more detailed categorization of failure modes.
The LLM planning failure modes we observed can be classified into the following three categories:
(1) Constraint Violation: The model fails to satisfy one of the explicitly declared environment constraints. i.e., it ignores one of the conditions required to be able to do an action. Even if this happen only for one instance of an action and not for all instances of the same action. For example model wants to put block b1 on block b2 while block b2 is not empty.
(2) Failure to meet the goal: The model makes a plan that meets the environment constraints but does not reach the goal state at the end of the trajectory.
(3) Out-of-vocabulary actions: The model proposes actions which are not in the environment’s action space.
PDDL benchmarks Here we examine the failure modes for both the Blocksworld and Logistics domains and SFT, ICL setting. Due to space constraints we report Blocksworld SFT settings here and refer the reader to Appendix C for the rest of plots and analysis. First we look at the effect of distribution of number of blocks in the planning instance. We compare Figure 5(a) that includes successful eval cases of BlocksWorld 8-20 blocks, to all its eval data in Figure 5(b) respectively. We note that the model failure cases increases as the number of blocks in the problem increases. We also separate the reasons of failure in BlocksWorld 8-20 instances as seen in Figure 5(c) and Figure 10(d) respectively.Note that failure mode (3) does not happen for BlocksWorld benchmark in SFT or ICL settings and for Logistics in ICL setting, but for Logistics benchmark in SFT settings all three categories are present (see Appendix C for details).
(a) Successful cases
(b) All cases
(c) Constraints Not Met
(d) Goal not reached
Figure 5: In-domain failure analysis: Distribution of number of blocks in successful and failed cases and different failure reasons for BlocksWorld 8-20 blocks. As the number of blocks increases the number of successful cases decreases.
Next, We study the distribution of number of blocks in out-of-domain cases focusing on BlocksWorld 3-7 to 8-20 blocks generalization. Comparing the successful cases in Figure 6(a) to all the eval samples in Figure 5(b), we observe that the generalization performance of the model that is trained on 3-7 blocks drops as the number of eval blocks increases from 8 to 20.
To probe the models even further, we look into the category where the generated plan fails to satisfy one of the environment constraints. We study the step at which the plan fails to meet the constraint. Considering Figure 6(b) and comparing it to Figure 6(c), the model seems to have trouble generalizing from the beginning, having the majority of its failures concentrated on earlier steps. This trend stays true for Logistics benchmark as well and is due to the fact that models go deep before they go wide when they want to produce the plan output. In other terms, this gives the intuition that the model seems to get overwhelmed when encountering difficult unseen examples, leading to failure from the beginning by not satisfying the constraints.
In addition, looking at the granular level of which actions that cause the failure in Figure 6(b), we note that the model has learned a correlation between the step number and the action it chooses, leading to having stack and un-stack not having any mutual failure step. This is also observed in failure modes of BlocksWorld 3-7 both for ICL and SFT scenarios (see Figure 9 in Appendix C). Further investigation of training data confirms this observation.
(a) Success. OOD generalization
(b) Fail. OOD generalization
(c) All cases
Figure 6: BlocksWorld 3-7 to 8-20 OOD scenario (a) Success per number of blocks (b) investigation of failures in constraint violation per plan step, color coded by action where the constraint violation happens (c) distribution of all cases per plan steps.
(a) TravelPlan errors by length
(b) Calendar Plan errors by length
Figure 7: Error Analysis for various CoT reasoning strategies in the TravelPlan and CalendarPlan natural language tasks
Natural Language benchmarks Since TravelPlan and CalendarPlan benchmarks require plans that do not involve sequential execution of actions, the only failure mode we consider is failure mode (2) - failure to reach a goal state. All errors discussed in this subsection are in terms of whether or not the method under study succeeded in reaching the goal. Figure 7 shows the average error by CoT method as task complexity increases in the TravelPlan and CalendarPlan benchmarks. Figure 12 in Appendix C histograms of how many examples belong to each task length. Both domains have a higher distribution of "smaller" or shorter tasks, which impacts the distribution of error in Figure 7. It is interesting to note how the error is disproportionate between debate-as-reasoning and MCTS, ToTs, aligning with our findings in Section 3. As discussed in Section 3, methods such as MCTS and ToT seem to perform inconsistently at less-structured planning tasks. We can see this in how they perform relative to debate, as well as each other. Specifically, for the TravelPlan task, they fail at inconsistent rates relative to one another.
5 Conclusion
We examined planning capabilities of LLMs through benchmark development, generalization assessment and analysis of failure modes. Our observations for ICL setting has implications for the future development and training of LLMs, potentially informing strategies to enhance their capacity to process and leverage extended contextual information. It also points to the opportunity for followup work investigating ways to further improve the inference time reasoning procedures, and their scaling properties. Our investigation of plan generalization reveals three key findings: superiority of SFT, curriculum learning effectiveness and limitations of hard example training; suggesting that a balanced approach, incorporating both easy and hard examples, might be optimal for achieving well-rounded performance. Our analysis of failure modes point to future work on dataset design and reasoning procedures to directly address the discussed failure modes. These enhancements can unlock new levels of versatility and robustness in LLM-based planning systems, paving the way for their broader adoption in real-world applications.
@@ B.1 Dataset Creation
In these experiments BlocksWorld dataset for 3 to 7 blocks consists of 40000 samples.
Algorithm 1 Create BlocksWorld Dataset
function CreateDatasetBW($num\_blocks$, $n$)
     $dataset\leftarrow[]$triangleright Initialize an empty list
     for $problem\_id\leftarrow 1$ to $n$ do
         $b\leftarrow$ RandomUniform($3$, $num\_blocks$)
         $initStacks\leftarrow$ CreateStacks($b$)
         $goalStacks\leftarrow$ CreateStacks($b$)
         if $initStacks==goalStacks$ then
              continue $\triangleright$ Skip equal stacks.
         end if
         $problem\leftarrow$CreatePro($initStacks,goalStacks$)
         $plan\leftarrow$FastDownward($problem,domain$)
         $dataset\leftarrow dataset+[(problem,plan)]$
     end for
     return $dataset$
end function
@@ B.2 Mappings PDDL to Natural Language
drive-truck param2 param3 param4 param5 Drive truck param2
from param3 to param4 in param5.
@@
unlock param2 param3 param4 param5 Unlock param2 at param3
using param4,
which has param5.